LINE MIXING COEFFICIENTS IN THE ν_{3} BAND OF ^{12}CH_{4} AND ^{13}CH_{4}

\underline{D. CHRIS BENNER}, V. MALATHY DEVI, Department of Physics, College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; MARY ANN H. SMITH, CURTIS P. RINSLAND, Atmospheric Sciences Division, MS 401A, NASA Langley Research Center, Hampton, VA 23681-2199; GUY GUELACHVILI, Laboratoire de Physique Moleculaire et Applications, CNRS, Universite Paris Sud, Batiment 350, 91405 ORSAY-Cedex, France; LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

Over 40 high-resolution absorption spectra of methane as previously described^{a}^{b} were recorded. Included in the spectra recorded at Kitt Peak were room temperature spectra of methane highly enriched in ^{13}CH_{4} broadened by air and having a 5 cm path length. All of these spectra were simultaneously fitted^{c} in order to obtain spectral line parameters in the ν_{3} spectral region of ^{12}CH_{4} and ^{13}CH_{4}. In order to obtain reasonable fits, it was necessary to include line mixing^{d} within the rotational manifolds. Line mixing was only required between spectral lines of the same species of methane. All of the P branch has been measured and measurement of the R branch is in progress. Mixing coefficients for the two isotopes and the two branches will be compared. It was also possible in a few cases to measure the line mixing coefficients between pairs of spectral lines belonging to the ν_{2}+ν_{4} band.

aV. Malathy Devi, D. Chris Benner, M. A. H. Smith, C. P. Rinsland, G. Guelachvili and L. R. Brown, Self- and Air-Broadening and Shift Coefficients of CH_{4} lines in the 3 μm Region, this conference (1998).

bV. Malathy Devi, D. Chris Benner, Mary Ann H. Smith, Curtis P. Rinsland, Guy Guelachvili and Linda R. Brown, Temperature Dependence Of Air-Broadening and Shift Coefficients in the ν_{3} Band of ^{12}CH_{4}, this conference (1998).

cD. C. Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT 53, 705-721 (1995).

dD. C. Benner, A Multispectrum Nonlinear Least Squares Fitting Technique: Inclusion of Line Mixing, this conference (1998).